Fermi National Laboratory
 

Operation Summary for November 15, 2010 - November 17, 2010

 

Wednesday Nov. 10

The day shift began with Operations monitoring Tevatron (TeV) store 8255 (89.67E30) and the Recycler stash (69.1E10), with the Antiproton Source (Pbar) stacking (11.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-992 experiments taking beam (FTBF T-1004 continues to take beam as a secondary user), and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam.

During the day shift, Operations terminated store 8255 at 3:05 PM and put the TeV into HEP shot setup. Recycler personnel investigated an Electron Cooling (ECool) BPM issue.

During the evening shift, Operations established store 8256 at 4:30 PM with an average initial luminosity of (AIL) 229E30. Water personnel added water to the MI LCW system following shot setup.

Thursday Nov. 11

The midnight shift began with Operations monitoring store 8256 (79.97E30) and the stash (211.8E10), with Pbar stacking (10.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that the MI south anode power supply tripped off; operators reset the supply. Operations terminated store 8256 at 7 AM and put the TeV into shot setup.

During the day shift, Operations established store 8257 at 8:15 AM with an AIL of 344E30. ECool personnel accessed MI-31. Operations reported that a Pbar Accumulator quadrupole power supply (A:QSF) developed a LCW leak and required about 30 minutes to repair. Operations transferred antiprotons to the Recycler until the stash reached 150E10.

During the evening shift, the ECool access continued. Operations reported that the MI-20 lower bus power supply tripped off on a filter choke overcurrent indication and was removed from the power supply turn-on order as requested by EE Support personnel. The Recycler maintained the 150E10 stash.

Friday Nov. 12

The midnight shift began with Operations monitoring store 8257 (69.91E30) and the stash (153.1E10), with Pbar stacking (158.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the ECool personnel completed their work on the Pelletron. Operations resumed transferring antiprotons from Pbar to the Recycler. Operations terminated store 8257 at 4:45 AM and put the TeV into shot setup. Unfortunately, D0 detector problems required Operations to halt shot setup. Operators prepared D0 for access. Operations reported problems with Pbar shot setup sequencer controls.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. CDF has a detector problem that is keeping them from taking data. The Run Coordinator will allow them to access on Saturday, 11/13/10, to repair their detector. This might take from 10 hours to two days. During the CDF access, the Run Coordinator will allow ECool personnel to work on the Pelletron, for Water personnel to repair a LCW leak in MI, and allow other machines to conduct minor repairs or studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/17/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory